THE BRAUER TREE OF THE PRINCIPAL 19-BLOCK OF THE SPORADIC SIMPLE THOMPSON GROUP

GERHARD HISS, KLAUS LUX, AND JÜRGEN MÜLLER

Abstract. This paper completes the construction of the Brauer tree of the sporadic simple Thompson group in characteristic 19. Our main computational tool to arrive at this result is a new parallel implementation of the DirectCondense method.